# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

### v2

@@ IV.4 Routing methods:
There are two routing algorithm choices. The first one, introduced in [51, 44], is a shuttle-based SWAP routing algorithm. In particular, this algorithm was tailored around the unique constraints of the crossbar architecture [38], which necessitated the maintenance of the checkerboard physical qubit pattern to achieve a time-efficient compilation process. One of the advantages of this algorithm is, in fact, the maintenance of the checkerboard pattern, which in turn keeps the crosstalk effect low and enables the Semi-Global rotations scheme exclusively. The second routing option is called beSnake [50], and it is capable of freely shuttling qubits around any topology and in any direction, handling complex routing scenarios involving parallelized gates.
@@ IV.5 SWAP replacement:
This binary variable concerns an optional functionality of the beSnake routing algorithm, which can replace a sequence of shuttles with a SWAP operation under certain conditions [50]. These conditions are satisfied when the accumulated fidelity of a parallelized shuttle sequence exceeds the fidelity of only executing a SWAP on that particular location. This variable is useful for creating insights for architectures that support SWAPs compared to others that do not support them.
@@ IV.6 Design interdependencies and verification
All these variable values act as input into the upgraded SpinQ housing its new configurable compiler, making it the first compilation and DSE framework for spin-qubit architectures. After this upgrade, we are able to provide SpinQ with multiple combinations of variable values, compile each input circuit iteratively for all valid architectures, and calculate and store the same performance metrics used in [51]. These are gate overhead, circuit depth overhead, and Estimated Success Probability (ESP) of the compiled quantum circuit(s).
@@
Another significant component of SpinQ is the compiler verification tool [51]. This tool is crucial, given the absence of real devices for testing. The verification functions have been enhanced to handle all valid architectures via our two-step method, which thoroughly scrutinizes circuits at each compilation stage [51].
@@ V Figure of Merit
To estimate the performance of any quantum processor, a prevalent method when the device is not physically available is through the Estimated Success Probability (ESP) [57, 47] extracted from the compiled circuit. This approach is by far the most efficient compared to Schrodinger-Feynman, tensor networks, or noise model simulations as all become exponentially slower with larger circuits and/or higher qubit connections on the device [51]. The choice of the figure of merit is a crucial step to ensure fair results, as every design variable needs to be directly linked to the resulting value, along with their trade-offs. This will prevent ArtA (introduced later in Section VI) from maximizing some variable(s) and, by doing so, unfairly favoring specific architectures. For example, it will not be fair to be able to maximize the parallelization degree of all gates without causing some negative impact due to increased crosstalk. Such a consideration also expresses more accurately design decisions in real experimental quantum processors.
In this work, we start from the ESP definition of [51] given in Equation 1 and expand it to reflect all previous considerations, given in Equation 2, with now three components in total: a gate operations term, a crosstalk term, and a decoherence term.
@@
where $k$ represents the $k$th time step and $i$ the $i$th gate in the $k$th time step.
$ESP=(\prod_{k}\prod_{i}{F_{i,k}})\cdot(\prod_{k}\prod_{i,j}C_{i,j,k})\cdot(1-e^{(-t/T_{2})})^{N}$ (2)
where $k$ represents the $k$th time step, $i$ the $i$th gate in the $k$th time step and $j$ represents the $j$th gate that inflicts crosstalk with the $i$th gate. $F$ is the fidelity and $C$ the crosstalk of the corresponding gates. The last term, adapted from [30], introduces the decoherence-induced errors, which represent the probability of all qubits staying coherent during the execution of the circuit. $T_{2}$ is the decoherence time, $t$ the duration time of the circuit, and $N$ the total number of qubits. We will discuss each of the terms in detail in the next three subsections.
@@ V.1 Gate operations term
The choice of typical operational fidelities of each gate, coined by the $F_{i,k}$ term, plays a large role in the probability of success of quantum circuits. In our analysis, we consider these fidelities to be constants within the design space rather than variables. This is because determining the optimal fidelity value for enhancing the ESP is straightforward; higher operational fidelity invariably leads to improved performance. Since we are conducting an exploration, we take a highly optimistic approach to the state-of-the-art [53, 75, 48] values for each gate type [^1] Another reason we do so is to avoid numeric underflows where there is a loss of accuracy in numerical calculations if ESP becomes smaller than the smallest positive representable value in a programming language’s floating-point arithmetic.:
@@ V.2 Crosstalk term
Crosstalk has been sparsely studied for spin qubits. This is because the crosstalk effects when executing multiple gates in parallel in close proximity are still very complex to model accurately [66, 29]. Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple forms [66]. For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures. For these reasons, we have devised a model to help us discover relative categorical differences between architectures, rather than accurate predictions of crosstalk effects. Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them. We further assume that crosstalk should depend on local hardware characteristics, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those. Below, we summarize our assumptions:
@@
A schematic overview of crosstalk effects is given in Figure 3. We see two two-qubit gates acting within the red ovals between Q-1 and Q-2, and Q-3 and Q-4, respectively. In this hypothetical square grid with one diagonal topology, crosstalk occurs in all three edges marked with a flash, as they directly connect the qubits of these two operations.
Figure 3: Schematic overview of cross-talk when executing two-qubit gates. Dashed black lines indicate a direct connection between two qubits, the red ovals indicate an ongoing two-qubit gate, and the blue flash icons indicate contributing crosstalk effects.
@@ V.3 Decoherence term
The last term of Equation 2 represents the decoherence of quantum information, which creates a strong trade-off between parallelization and crosstalk effects. Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
@@
where $D_{i,k}$ is the duration of the $i$th operation in the $k$th time step. In Table 2, we present the gate duration and decoherence time used. The motivation behind this selection of durations, once again, is for purposes of architectural comparisons. We took reference for our two-qubit gate (i.e., $\sqrt{SWAP}$) from [48, 75], for the $Shuttle$ from [30] which is half the shuttle-based Z rotations [51], for the pulsed-based single-qubit gates from [53], and the $T_{2}$ from [31].
@@
VI ArtA (Artificial Architect)
@@
From the aforementioned transformation of the SpinQ compilation framework to a Design Space Exploration (DSE) framework, we are able to brute force through all valid combinations of the input variable values. However, as we discussed, this can be a time-consuming process, especially when the defined design space is large. Therefore, a specialized tool is needed to automate the DSE process by traversing the space in a clever and faster way. Ultimately, this will result in fewer designs being tried until desired ESP values or other metrics are reached.
To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers. Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit and universally.
The abstracted ArtA process is depicted in Figure 4. The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Section IV, the optimization method(s) to traverse the space, and an (optional) termination condition. Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself. This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (ESP in this work). Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored. In the following sections, we will describe the TC and ArtA metrics in more detail.
@@
To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities. In the first one, we answer the third research question by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit. Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP. Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2 evaluating ArtA’s overall performance. In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.
VI.1 Termination Condition (TC)
During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used. Since we first obtain the ESP values for all architectures during the third question, the highest ESP is known and can work as a TC. Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC. More specifically, our analysis revealed that over 95% of processes are completed earlier, while a minority extends significantly longer. Given that our evaluation focuses on runtime, omitting these prolonged instances won’t affect the overall assessment, as their inclusion would lead to unfavorable evaluations regardless.
@@ VI.1 Termination Condition (TC)
For the second use of ArtA, which answers the fourth research question, we investigate the returned architectural designs by using the chosen optimization method derived from answering the third question. We find in practice that the corresponding chosen optimization algorithm did not exceed trying 23.4% of the total $29,312$ architectures in the worst case before finding the one with the highest ESP.
@@ VI.2 Metrics Evaluating ArtA
To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics. The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach of SpinQ. Otherwise, ArtA will be of no practical interest. We thus use the relative time-to-solution as a metric in Equation 6. The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute force the entire design space.
@@
When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method slowly becomes less important within the total runtime. Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance. Thus, a relative calls-to-solution metric, in Equation 7, is also a strong performance predictor for ArtA, especially for larger circuits. With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture. In case the optimization method attempts the same architecture twice or more, the compilation will be skipped since the results are already stored from the first call. Finally, the denominator represents the brute force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
@@ VII Results and Evaluation
In this section, we will attempt to completely answer the third and fourth research questions. In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space. Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions. Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TC. Besides the mean performance, we will also report the worst performing results for reference.
@@ VII.1 Comparison of optimization methods
Table 3: Relative time-to-solution metric (the lower, the better) for finding the architecture with the highest ESP from the design space. A minus symbolizes that there was at least one run (out of the ten) in which the top ESP value was not reached in time (TC = 40 minutes). These results are colored red, as well as runs where the time taken exceeds the brute force time (values over 100). The green cells indicate the top five configurations of optimization methods for a given quantum circuit. The mean and worst performing configurations from a sample of ten runs are shown for each quantum circuit (block of rows). QFT: Quantum Fourier Transformation, cA: Cuccaro Adder, vA: vbe Adder, QV: Quantum Volume, G: Grover’s, RC: Random Circuit, RS: Random Sampling, SA: Simulated Annealing (Top: Start Temperature, bottom: Step Size), BO: Bayesian Optimization (Top: Acquisition function, bottom: Kernel), GA: Genetic Algorithm (Top: Population Size, bottom: Mutation), ACO: Ant Colony Optimization (Top: Population Size, bottom: Exploitation Rate).
@@
In this first part of our results, we will address the third research question stated in Section III, which is about finding the best performant optimization method according to our ArtA metrics defined in Section VI.2. The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3. This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits. As explained previously, the denominator of Equation 6 is obtained beforehand by measuring the time taken by the brute force approach. We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100). For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute forcing all architecture, and 100 means that it takes the same time.
@@
Multiple optimization methods greatly outperform the random sampling (RS column), as expected. The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute force time to find the architecture with the highest ESP.
@@
We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%. Overall, different optimization methods clearly show varied performance across their respective columns. However, the performance differences among various hyper-parameter settings within a single optimization method are less distinct. In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA. More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%. Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results across the ten runs. To further support our findings, next, we will assess the relative calls-to-solution metric.
@@
The best optimization method only takes 1% of the total compiler calls, on average, for QV8 and QV10 with ACO [50, 0.1]. We calculated based on Table 4 that, on average, ArtA’s optimization method configurations can decrease the compiler calls by 79.9%. We can observe, once again, that BO, followed by GA, obtained the best performance compared to other optimization methods. Lastly, in 91.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, showing results consistency again.
@@
From the different variants of the GA and BO used, we focus on GA [100, 0.2] and BO [EI, MA], as they have the best performance and did not reach our TC in any of the ten runs. By comparing the two tables, we can observe that BO [EI, MA] needs marginally more time to evaluate fewer architectures compared to GA [100, 0.2]. Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, MA] compared to GA [100, 0.2]. This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Section X.2). However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits. Based on their small relative time-to-solution difference, we can predict that BO [EI, MA] will outperform GA [100, 0.2] for larger circuits as the latter has, on average, 58% higher relative calls-to-solution. Therefore, we conclude that BO [EI, MA] is a better candidate for automating the DSE process for large-scale circuits, as it can traverse the design space faster.

### v3

@@ IV.4 Routing methods via shuttling:
There are two available algorithmic options. The first one, introduced in [56, 48], is a shuttle-based SWAP routing algorithm. In particular, this algorithm was tailored around the unique constraints of the crossbar architecture [41], which necessitated the maintenance of the checkerboard physical qubit pattern to achieve a time-efficient compilation process. One of the advantages of this algorithm is, in fact, the maintenance of the checkerboard pattern, which in turn keeps the crosstalk interference low and exclusively enables the Semi-Global rotations scheme. The second routing option is called beSnake [55], and it is capable of freely shuttling qubits around any topology and in any direction, handling complex routing scenarios involving parallelized gates.
@@ IV.5 SWAP replacement:
This binary variable concerns an optional functionality of the beSnake routing algorithm, which can replace a sequence of shuttles with a SWAP gate under certain conditions [55]. These conditions are satisfied when the accumulated fidelity of a parallelized shuttle sequence exceeds the fidelity of only executing a SWAP on that particular location. This variable is useful for creating insights for architectures that support SWAPs compared to others that do not support them.
@@ IV.6 Design interdependencies and verification
All these variable values act as input into the upgraded SpinQ housing its new configurable compiler, making it the first compilation and DSE framework for spin-qubit architectures. After this upgrade, we are able to provide SpinQ with multiple combinations of variable values, compile each input circuit iteratively for all valid architectures, and calculate and store multiple performance metrics, such as the ones used in [56]. These are gate overhead, circuit depth overhead, and ESP of the compiled quantum circuit(s).
@@
Another significant component of SpinQ is the compiler verification tool [56]. This tool is crucial, given the absence of real devices for testing. For this work, the verification functions have been enhanced to handle all valid architectures via our two-step method [56], which thoroughly scrutinizes circuits at each compilation stage.
@@ V Figure of merit
The selection of an appropriate figure of merit is a critical step to ensure fair and objective evaluation of quantum processor architectures. It must reflect how specific design features influence circuit performance, allowing each architecture to be characterized based solely on its inherent trade-offs and capabilities. To achieve this, each design variable must contribute directly to the final metric, such that the interplay between performance gains and associated costs (e.g., crosstalk or decoherence) is properly captured. This prevents the optimization tool ArtA (introduced in Section VI) from disproportionately favoring certain architectures by, for instance, maximizing the degree of gate parallelization without accounting for the negative effects it may introduce, such as increased crosstalk. To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk. Such considerations also expresses more accurately design decisions in real experimental quantum processors.
When a quantum processor is not physically available, a common method for estimating its performance is to compute the Estimated Success Probability (ESP) [63, 51, 57], derived from the compiled quantum circuit. This method is significantly more computationally efficient than alternatives such as Schrödinger-Feynman simulation, tensor network contraction, or noise model simulations, all of which scale poorly with increasing circuit size and qubit connectivity [56, 57]. In this work, we build upon the ESP definition introduced in [56], shown in Equation 1, and extend it to incorporate all previously discussed architectural considerations. The revised figure of merit, presented in Equation 2, now consists of three components: a gate operations term, a crosstalk term, and a decoherence term.
@@
where $k$ represents the $k$th time step and $i$ the $i$th gate in the $k$th time step. $F$ is the fidelity of the corresponding gates.
$ESP=(\prod_{k}\prod_{i}{F_{i,k}})\cdot(\prod_{k}\prod_{i,j}C_{i,j,k})\cdot{\color[rgb]{0,0,0}(e^{(-t/T_{2}^{*})})^{N}}$ (2)
where $C$ denotes the crosstalk between the $i$th and $j$th gates during the $k$th time step. The last term, adapted from [30], introduces the decoherence-induced errors, which represent the probability of all qubits staying coherent during the execution of the circuit’s execution. $T_{2}^{*}$ is the decoherence time, $t$ the total duration time of the circuit, and $N$ the total number of qubits. We will discuss each of the terms in detail in the next three subsections.
@@ V.1 Gate operations term
The selection of typical operational fidelities for each gate, denoted by the $F_{i,k}$ term, plays a significant role in determining the success probability of quantum circuits. In our analysis, we consider these fidelities to be constants within the design space, rather than variables. This is because determining the optimal fidelity value for increasing the figure of merit (i.e., ESP) is straightforward; higher operational fidelity invariably leads to improved performance. Since we are conducting an exploration, we take a highly optimistic approach to the state-of-the-art [59, 86, 52] values for each gate type [^1] Another reason we do so is to avoid numeric underflows where there is a loss of accuracy in numerical calculations if ESP becomes smaller than the smallest positive representable value in a programming language’s floating-point arithmetic.:
@@ V.2 Crosstalk term
Crosstalk in spin qubit systems has received limited attention in the literature, primarily due to the complexity of accurately modeling the effects that arise when multiple gates are executed simultaneously in close physical proximity [75, 29, 37]. Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple occurrences [75]. For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures. For these reasons, we have devised a model to help us discover relative categorical differences between architectures, rather than accurate predictions of crosstalk effects. Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them. We further assume that crosstalk should depend on local hardware characteristics, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those. Below, we summarize our assumptions:
@@
A schematic overview of crosstalk effects is presented in Figure 3. The diagram illustrates two two-qubit gates, represented by red ovals, acting between qubits Q-1 and Q-2, and Q-3 and Q-4, respectively. In this hypothetical architecture with degree = 4 connectivity, crosstalk arises along the three edges marked with a lightning symbol, as the involved qubits are directly connected to one another.
Figure 3: Schematic overview of crosstalk during the execution of two two-qubit gates. Dashed black lines represent direct qubit connections, red ovals indicate active two-qubit gates, and blue lightning icons denote edges where crosstalk effects occur.
@@ V.3 Decoherence term
The last term of Equation 2 represents the decoherence of quantum information, which creates a strong architectural trade-off between parallelization and crosstalk effects. Since $N$ and $T_{2}^{*}$ are given, we need to define the duration time of the circuit $t$:
@@
where $D_{i,k}$ is the duration of the $i$th operation in the $k$th time step. In Table 2, we present the gate duration and decoherence time used. The motivation behind this selection of durations, once again, is for purposes of architectural comparisons. We took reference for our two-qubit gate (i.e., $\sqrt{SWAP}$) from [52, 86], for the $shuttle$ from [30] which is half the shuttle-based Z rotations [56], for the pulse-based single-qubit gates from [59], and for $T_{2}^{*}$ from [31].
@@
VI ArtAArtificial Architect
From the aforementioned transformation of the SpinQ compilation framework to a fully functional DSE framework, we are able to brute-force through all valid combinations of the input variable values. However, as we discussed, this can be a time-consuming process, especially when the defined design space is large. Therefore, a specialized tool is needed to automate the DSE process by traversing the space in a clever and faster way. As a result, fewer design iterations will be tried to achieve the same ESP (range) or other specified performance metrics.
@@ VI ArtAArtificial Architect
To do that, we introduce ArtA, a tool integrated into the SpinQ framework that incorporates multiple optimization techniques. Its purpose is twofold: (a) to determine which optimization method is most effective at identifying architectures that achieve the desired ESP for each quantum circuit, and (b) to identify the architecture that yields the highest ESP both on a per-circuit basis and across all circuits collectively (i.e., a universal architecture).
The abstracted ArtA process is depicted in Figure 4. The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Section IV, the optimization method(s) to traverse the space, and an (optional) termination condition. Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture in the form of input variable values from the available “population” of architectures – a subset of design space selected by the optimization algorithm itself. This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired range/value of a specific metric is reached (ESP in this work). Upon termination, the resulting architecture, its ESP calculation, and ArtA metrics are stored. In the following sections, we will describe the TCs and ArtA metrics in more detail.
@@
To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities. With the first one, we will answer the third research question, in Section VII.1, by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit. Therefore, first we need to compile each circuit for all $29,312$ architectures and store their ESP for later use. Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2, evaluating ArtA’s overall performance. With the second capability of ArtA, we will only use the best-performing optimization method, but expand our circuit selection to concretely answer the last research question, in Section VII.2.
VI.1 Termination condition
During each optimization cycle, a specific TC is checked to reduce further computational time based on which way, (a) or (b), ArtA is used. Since we first obtain the ESP values for all architectures during the third question, as mentioned before, the highest ESP is known and can work as a TC. Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC. More specifically, our analysis revealed that over 95% of processes are completed earlier, while a minority extends significantly longer. Given that our evaluation focuses on runtime, omitting these prolonged instances won’t affect the overall assessment, as their inclusion would lead to unfavorable evaluations regardless.
@@ VI.1 Termination condition
For the second use of ArtA, which answers the last research question, we investigate the returned architectural designs by using the best optimization method derived from answering the third question. As we will see later in Section VII.1, we find that the corresponding chosen optimization algorithm did not exceed 23.4% of the total $29,312$ architectures in the worst case before finding the one with the highest ESP.
@@ VI.2 Metrics evaluating ArtA
To determine the performance of ArtA, in the context of replying to the third research question, we need to define some relevant metrics. The optimization methods that will eventually determine ArtA’s performance should return a solution faster compared to searching with the native brute-force approach of SpinQ. Otherwise, ArtA will be of no practical interest. We thus use the relative time-to-solution as a metric in Equation 6. The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute-force the entire design space.
@@
When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method becomes relatively less significant in the total runtime. This happens because, as circuits grow in size, the compilation time per architecture dominates the total runtime, while the time taken by the optimization method remains approximately constant regardless of circuit size. Therefore, time-to-solution is not sufficient in evaluating ArtA’s performance, as it becomes increasingly influenced by compilation time, rather than the efficiency of the optimization method itself. To address this, we also introduce a relative calls-to-solution metric in Equation 7. With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture. In case the optimization method attempts the same architecture twice or more, the circuit compilation will be skipped since the results are already stored from the first call. Finally, the denominator represents the brute-force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
@@ VII Results and evaluation
In this section, we will attempt to completely answer the third and fourth research questions. In both questions, we will focus on finding the architecture that obtains the highest ESP among all within the design space. Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions. Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TCs. Besides the mean performance, we will also report the worst performing results for reference.
@@ VII.1 Comparison of optimization methods
Table 3: Relative time-to-solution metric (the lower, the better) for finding the architecture with the highest ESP from the design space. A minus symbolizes that there was at least one run (out of the ten) in which the top ESP value was not reached in time (TC = 40 minutes). These results are colored red, as well as runs where the time taken exceeds the brute-force time (values over 100). The green cells indicate the top five configurations of optimization methods for a given quantum circuit. The mean and worst performing configurations from a sample of ten runs are shown for each quantum circuit (block of rows). QFT: Quantum Fourier Transformation, cA: Cuccaro Adder, vA: vbe Adder, QV: Quantum Volume, G: Grover’s, RC: Random Circuit, RS: Random Sampling, SA: Simulated Annealing (Top: Start Temperature, bottom: Step Size), BO: Bayesian Optimization (Top: Acquisition function, bottom: Kernel), GA: Genetic Algorithm (Top: Population Size, bottom: Mutation), ACO: Ant Colony Optimization (Top: Population Size, bottom: Exploitation Rate).
@@
In this first part of our results, we will address the third research question stated in Section III, which is about finding the best performant optimization method according to our ArtA metrics defined in Section VI.2. The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3. This table also provides insights into matching different optimization method configurations (described in Appendix X.2) with specific quantum circuits. We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100). For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute-forcing all architecture, and 100 means that it takes the same time.
@@
Multiple optimization methods greatly outperform the random sampling (RS column), as expected. The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute-force time to find the architecture with the highest ESP.
@@
We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%. Overall, different optimization methods clearly show varied performance across their respective columns. However, the performance differences among various hyper-parameter settings within a single optimization method are less distinct. In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA. More specifically, GA, with a population size of 100 and a mutation probability of 0.2, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, Ma] coming close at 12.1%. Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results across the ten runs. To further support our findings, next, we will assess the relative calls-to-solution metric.
@@
The best optimization method only takes 1% of the total compiler calls, on average, for QV8 and QV10 with ACO [50, 0.1]. We calculated based on Table 4 that, on average, ArtA’s optimization method configurations can decrease the compiler calls by 79.9%. We can observe, once more, that BO, followed by GA, obtained the best performance compared to other optimization methods. Lastly, in 91.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, showing results consistency again.
@@
From the different variants of the GA and BO used, we focus on GA [100, 0.2] and BO [EI, Ma], as they have the best performance and did not reach our TC in any of the ten runs. By comparing the two tables, we can observe that BO [EI, Ma] needs marginally more time to evaluate fewer architectures compared to GA [100, 0.2]. Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, Ma] compared to GA [100, 0.2]. This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Appendix X.2). However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits. Based on their small relative time-to-solution difference, we can predict that BO [EI, Ma] will outperform GA [100, 0.2] for larger circuits as the latter has, on average, 58% higher relative calls-to-solution. Therefore, we conclude that BO [EI, Ma] is a better candidate for automating the DSE process for large-scale circuits, as it can traverse the design space faster.